Of 150 workers in a town surveyed, 65 work in the mall and 30 work in the factory. What is the probability that a randomly selected worker works in either the mall or the factory? What is the probability that they work in neither place? What assumptions did you make?
Answer #1

n(mall) = 65
n(factory) = 30
n(mall or factor)' = 150 - 95 = 55

P(mall or factory) = (65 + 30)/150 = 95/150 = 0.6333
Here the assumption made is that no worker is working at both mall and factory.

P(neither place) = 55/150 = 0.3667